# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

### v2

@@ Abstract
Randomly initialized dense networks contain subnetworks that achieve high accuracy without weight learning—strong lottery tickets (SLTs). Recently, Gadhikar et al. [10] demonstrated that SLTs can also be found within a randomly pruned source network, thus reducing the SLT search space. However, this limits the search to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity. This paper proposes a method that reduces the SLT search space by an arbitrary ratio independent of the desired SLT sparsity. A random subset of the initial weights is excluded from the search space by freezing it—i.e., by either permanently pruning them or locking them as a fixed part of the SLT. In addition to reducing search space, the proposed random freezing can also provide the benefit of reducing the model size for inference. Furthermore, experimental results show that the proposed method finds SLTs with better accuracy-to-model size trade-off than the SLTs obtained from dense or randomly pruned source networks. In particular, the SLTs found in Frozen ResNets on image classification using ImageNet significantly improve the accuracy-to-search space and accuracy-to-model size trade-offs over SLTs within dense (non-freezing) or sparse (non-locking) random networks.
@@ 1 Introduction
The Strong Lottery Ticket Hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks [41, 30, 25]. The existence of such subnetworks that do not require weight training, illustrated in Figure 2(left), has been demonstrated experimentally [41, 30, 17, 21, 39] and proven theoretically for dense source networks [25, 27, 28, 6, 5, 1, 3, 8, 10].
@@
SLTs offer a particularly advantageous opportunity for specialized inference hardware [14] since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the sparse binary pruning masks—supermasks—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads. Furthermore, the binary nature of both the supermasks and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al. [14]. On the other hand, the existing training algorithms for finding SLTs suffer from the same high computational cost as weight optimization, since they also employ backpropagation to update one parameter per neural connection [30, 42, 6, 33].
@@
Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the optimization targets for finding an SLT.
@@
Figure 2: Freezing ($\bullet$) improves the accuracy-to-search space trade-off over pruning only (non-locking) ($\bullet$) or non-freezing ($\bullet$).
Recently, Gadhikar et al. [10] demonstrated that this steep training cost can be alleviated by randomly pruning the source network at initialization, reducing the number of edges to be optimized (see Figure 2, center). However, by its nature, this method can only search SLTs in relatively high sparsity region, which also includes highly sparse networks with terrible accuracy such as layer collapse [11, 34]. Therefore, it would be desirable if we could reduce the search space while avoiding too sparse networks.
@@
This paper introduces a novel method that reduces the search space by an arbitrary ratio independent of the desired SLT sparsity: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in  Figure 2(right). Both the pruned and the locked parameters—the frozen parameters—are excluded from the search space, therefore reducing the training cost. Far from negatively impacting performance, this cost reduction proves efficient: the SLTs found within this partial search space are more accurate than those resulting from conventional methods, as shown in Figure 2. Moreover, resulting from the fact that this method can also be exploited to reduce model size in inference, our experiments show that SLTs found within frozen networks can achieve a better accuracy-to-model size trade-off than SLTs obtained from dense [30, 33] or sparse [10] random networks.
The contributions of this paper can be summarized as follows:
  • We propose a novel method that reduces the SLT search space with any desired sparsity by random partial freezing (pruning and locking) of the source random network. The soundness of our method is theoretically supported by our existence theorem of SLTs in such a randomly reduced search space, based on subset-sum approximation with randomly frozen variables.
@@
  • We experimentally validated our method in three scenarios with highly accurate SLTs in low, medium, and high SLT sparsity regions, which reveal that parameter freezing consistently reduces the size of search spaces. Even with randomly frozen parameters, our method finds highly accurate SLTs that cannot be found within dense networks in some desired sparsities.
@@
  • Furthermore, the experimental results show that SLTs found in the frozen network achieve better accuracy-to-search space and accuracy-to-model size trade-offs than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
@@
As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major source of energy and time consumption [15]—can be drastically reduced for energy-efficient inference acceleration [14]. These contributions could potentially lead to attracting attention to SLTs also as an approach for efficient training, acquiring high-quality subnetworks with reduced training costs.
@@ 2 Preliminaries
After outlining the background on strong lottery tickets (SLTs) and their search algorithms, this section recapitulates the subset-sum approximation technique for the existence proof of SLTs [28, 5, 2, 1] and its extension to sparse random networks [10], which Section 3 expands to the existence of SLTs within frozen networks.
@@ 2.1 Strong Lottery Tickets in Dense Networks
SLTs [41, 30, 25] are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training. Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed [14]. Instead of updating weights, SLT search algorithms for deep neural networks [30, 42, 33] use backpropagation to update weight scores, which are then used to generate the supermask. For example, the Edge-Popup algorithm [30, 33] used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores. The need to keep scores in addition to the weights and the score sorting operation necessary for generating the supermask make the computational cost of Edge-Popup slightly higher than conventional weight training.
Figure 3: We extend the SLT existence proof by Burkholz [1] (left) to frozen source networks (right) by solving the subset-sum approximation with some source parameters that are either unavailable ($w^{\prime}_{2}$) or must be used for the subset-sum ($w^{\prime}_{0}$).
2.2 SLT Existence via Subset-Sum Approximation
Given a set of random variables and a target value, the subset-sum problem consists on finding a subset whose total value approximates the target. Lueker [24] showed that such subset exists with high probability if the number of random variables is sufficiently large:
Lemma 2.1 (Subset-Sum Approximation [24]).
@@ 2.2 SLT Existence via Subset-Sum Approximation
Based on Lemma 2.1, which was first introduced into the SLT context by Pensia et al. [28], previous works [1, 2, 5, 28] showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network. In particular, Burkholz [1] proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$. As exemplified in Figure 3 (left), to approximate a weight $w_{T}$ of the target network, the SLT uses weights $w_{0},w_{1}$, and $w_{2}$ from the source network to construct the subset-sum. Therefore, the source network is required to be sufficiently wider than the target network, as stated in Lemma 2.1.
@@ 2.3 Strong Lottery Tickets in Sparse Networks
Recently, Gadhikar et al. [10] revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center). In addition to their theoretical contributions, their approach offered a practical advantage: since there are less parameters available from the beginning, the search space is reduced in proportion to the initial pruning ratio. This search space reduction can be exploited to reduce the computational cost of training.
2.4 SLT Existence in Sparse Networks
To prove the existence of SLTs within sparse networks, Gadhikar et al. [10] extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
Lemma 2.2 (Subset-Sum Approximation in Sparse Networks [10]).
@@ 2.4 SLT Existence in Sparse Networks
By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz [1], Gadhikar et al. [10] proved the SLT existence in sparse networks as the following theorem:
Theorem 2.3 (SLT Existence in Sparse Networks [10]).
@@ 3 Strong Lottery Tickets in Frozen Networks
Although pruning at initialization provides a search space reduction, it also imposes a limitation on the sparsity range, as it can only search for SLTs with a sparsity higher than the initial pruning ratio. For example, an initial pruning ratio of $90\%$ vastly reduces the search space, but it limits the search to only SLTs sparser than $90\%$. Therefore, this reduction in search space is incompatible with exploring the whole sparsity range for optimal accuracy.
This paper proposes a novel method that reduces the search space regardless of the desired SLT sparsity to capitalize on the training cost reduction without sacrificing accuracy. We explore the existence and performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right). Since freezing is performed at initialization before the SLT search, frozen parameters do not have a score and are not eligible for exploration, thus reducing the search space. Locking allows to further reduce the parameter search without limiting the initial pruning ratio, reducing the search space even in the case of dense source networks.
This section first provides a theoretical analysis of the SLT existence in such frozen networks by developing a new extension of the subset-sum approximation introduced in Section 2. Then, based on this analysis, it proposes partial SLT search, a novel method for finding such SLTs with a reduced search space. In preparation for the evaluation experiments in Section 4, we also perform a preliminary investigation of the optimal settings of the proposed method.

### v3

# Partially Frozen Random Networks Contain Compact Strong Lottery Tickets

@@ Abstract
Randomly initialized dense networks contain subnetworks that achieve high accuracy without weight learning—strong lottery tickets (SLTs). Recently, Gadhikar et al. (2023) demonstrated that SLTs could also be found within a randomly pruned source network. This phenomenon can be exploited to further compress the small memory size required by SLTs. However, their method is limited to SLTs that are even sparser than the source, leading to worse accuracy due to unintentionally high sparsity. This paper proposes a method for reducing the SLT memory size without restricting the sparsity of the SLTs that can be found. A random subset of the initial weights is frozen by either permanently pruning them or locking them as a fixed part of the SLT, resulting in a smaller model size. Experimental results show that Edge-Popup (Ramanujan et al., 2020; Sreenivasan et al., 2022) finds SLTs with better accuracy-to-model size trade-off within frozen networks than within dense or randomly pruned source networks. In particular, freezing $70\%$ of a ResNet on ImageNet provides $\times3$ compression compared to the SLT found within a dense counterpart, raises accuracy by up to $14.12$ points compared to the SLT found within a randomly pruned counterpart, and offers a better accuracy-model size trade-off than both.
@@ 1 Introduction
The strong lottery ticket hypothesis (SLTH) conjectured the existence of subnetworks within a randomly weighted network—strong lottery tickets (SLTs)—that achieve comparable accuracy to trained dense networks (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020). The existence of such subnetworks that do not require weight training, illustrated in Figure 2 (left), has been demonstrated experimentally (Zhou et al., 2019; Ramanujan et al., 2020; Huang et al., 2022; López García-Arias et al., 2023; Yeo et al., 2023) and proven theoretically for dense source networks (Malach et al., 2020; Orseau et al., 2020; Pensia et al., 2020; Diffenderfer & Kailkhura, 2021; da Cunha et al., 2021; Burkholz, 2022a; Burkholz et al., 2022; Ferbach et al., 2023; Gadhikar et al., 2023). They appear in networks with excessive amounts of parameters, and their existence indicates the possibility of new optimization strategies for deep neural networks based on a network connectivity perspective.
@@
SLTs offer a particularly advantageous opportunity for specialized inference hardware (Hirose et al., 2022; Chen et al., 2022; 2023) since the random weights can be reconstructed from the seed (i.e., they do not need to be memorized), and the binary mask—supermask—can be greatly compressed with entropy coding, vastly reducing off-chip memory access and its associated overheads. Furthermore, the binary nature of both the supermask and the random weights can be exploited for multiplier-less execution, as demonstrated practically by Hirose et al. (2022).
@@
Figure 1: Freezing the source network by randomly pruning some parameters and locking others reduces the memorized supermask for finding an SLT.
@@
Figure 2: Freezing ($\bullet$) improves the accuracy-to-model size trade-off over pre-pruning only ($\bullet$) or non-freezing ($\bullet$).
Recently, Gadhikar et al. (2023) showed that accurate SLTs could be found even if the number of edges to be optimized was reduced by randomly pruning the source network at initialization (see Figure 2, center). Since the random pre-pruning mask can be reconstructed from the seed in the same way as the random weights, and thus there is no need to store the part of the supermask corresponding to pre-pruned weights, their approach can be exploited to further reduce the memory cost required by the SLT in specialized hardware. However, when aiming for high compression, their method can only search for SLTs in the relatively high sparsity region, and may even lead to layer collapse (Hayou et al., 2020; Tanaka et al., 2020). Furthermore, since the SLT sparsity regions where highly accurate SLTs exist depend on the dataset and network architecture, a search limited to sparse regions may fail to find accurate SLTs. For instance, work on SLTs within graph neural networks has shown that dense SLTs are more accurate than sparse ones in some settings (Huang et al., 2022; Yan et al., 2023). Therefore, it would be desirable to use a method that allows for increased randomness in SLTs for further model compression, but has the freedom to allocate it in the SLT sparsity regions that lead to more accurate tickets.
@@
This paper introduces such a novel method to reduce the memory cost of the optimized supermask without restricting the desired sparsity of SLTs to be searched for: in addition to random pruning at initialization, it also locks randomly chosen parameters at initialization to be a permanent part of the SLT (i.e., never pruned), as exemplified in Figure 2 (right). Both the randomly pruned and the locked parameters—the frozen parameters—are left completely random and can be regenerated from seed. The weights corresponding to the optimized supermask region (less than $50\%$ of the total) are also reconstructed from seed, and the supermask is binary and sparse, producing a highly compressible model. Far from negatively impacting performance, this cost reduction is efficient: as shown in Figure 2, the frozen SLTs achieve a higher accuracy than SLTs with a similar size resulting from conventional methods. The contributions of this paper are summarized as follows:
  • We propose a novel method that vastly reduces the number of parameters to be memorized for finding an accurate SLT by freezing (pruning and locking) the source random network.
@@
  • We experimentally validate our method in three scenarios corresponding to low, medium, and high optimal SLT sparsity regions, which reveal that parameter freezing consistently produces smaller yet accurate supermasks. Even with randomly frozen parameters, we find highly accurate SLTs that cannot be found within dense networks for some desired sparsities.
@@
  • Furthermore, the experimental results show that SLTs found in frozen networks achieve comparable or better accuracy-to-model size trade-off than SLTs found within dense (non-freezing) or sparse (non-locking) random networks.
@@
As mentioned above, SLTs are quite attractive for neural engine design, as they can vastly reduce the memory size for model storage, meaning that off-chip memory access—by far the major bottleneck of energy and time consumption (Horowitz, 2014)—can be drastically reduced for energy-efficient inference acceleration (Hirose et al., 2022). Our contributions have the potential to reduce off-chip memory access further and to make inference more energy-efficient than previous designs.
@@ 2 Preliminaries
This section outlines the background of strong lottery tickets (SLTs) within dense or sparse networks and SLT search algorithms.
@@ 2.1 Strong Lottery Tickets in Dense Networks
SLTs (Zhou et al., 2019; Ramanujan et al., 2020; Malach et al., 2020) are subnetworks within a randomly weighted neural network that achieve high accuracy without any weight training. Compared with learned dense weight models, SLTs can be reconstructed from a small amount of information: since the random weights can be regenerated from their seed, it is only necessary to store the binary supermask and the seed (Hirose et al., 2022). SLT search algorithms for deep neural networks (Zhou et al., 2019; Ramanujan et al., 2020; Zhou et al., 2021; Sreenivasan et al., 2022) find SLTs by updating weight scores, which are then used to generate the supermask, instead of updating weights. For example, the Edge-Popup algorithm (Ramanujan et al., 2020; Sreenivasan et al., 2022) used in this paper finds SLTs by applying a supermask generated from the connections with the top-$k\%$ scores.
SLT Existence via Subset-Sum Approximation
Given a set of random variables and a target value, the subset-sum problem consists of finding a subset whose total value approximates the target. Lueker (1998) showed that such a subset exists with high probability if the number of random variables is sufficiently large:
Lemma 2.1 (Subset-Sum Approximation (Lueker, 1998)).
@@
Based on this Lemma 2.1, which was first introduced into the SLT context by Pensia et al. (2020), previous works (Burkholz, 2022a; b; da Cunha et al., 2021; Pensia et al., 2020) showed that an SLT that approximates an arbitrary target network exists in a dense source network if it is logarithmically wider and constantly deeper than the target network. In particular, Burkholz (2022a) proved that a source network with depth $L\!+\!1$ and larger width than the target network contains an SLT that can approximate the target network with depth $L$.
@@ 2.2 Strong Lottery Tickets in Sparse Networks
Recently, Gadhikar et al. (2023) revealed that SLTs also exist within sparse source networks, i.e., random networks that have been randomly pruned at initialization (see Figure 2, center). They showed its existence experimentally as well as theoretically, as outlined later: an SLT that approximates a given target network exists with high probability in a sparse source network that is sufficiently wider and deeper than the target.
SLT Existence in Sparse Networks
To prove the existence of SLTs within sparse networks, Gadhikar et al. (2023) extended the subset-sum approximation (Lemma 2.1) to the situation where randomly chosen variables are permanently pruned at initialization:
Lemma 2.2 (Subset-Sum Approximation in Sparse Networks (Gadhikar et al., 2023)).
@@
By applying this extended lemma instead of Lemma 2.1 to the SLT existence theorem presented by Burkholz (2022a), Gadhikar et al. (2023) proved the SLT existence in sparse networks as following theorem:
Theorem 2.1 (SLT Existence in Sparse Networks (Gadhikar et al., 2023)).
@@ 3 Strong Lottery Tickets in Frozen Networks
The approach of finding SLTs within sparse networks by Gadhikar et al. (2023) offers a practical advantage: as the randomly pre-pruned parts can be reconstructed from the seed, this pre-pruning process can be exploited to reduce the memory requirement of the supermask.
@@
However, this approach also imposes the limitation that it can only search for SLTs with a sparsity higher than the pre-pruning ratio. For example, a random pre-pruning ratio of $90\%$ vastly reduces the memory size of the supermask, but it limits the search to only SLTs sparser than $90\%$. Therefore, their method is incompatible with exploring the whole SLT sparsity range for optimal accuracy.
This paper proposes a novel method that allows us to find SLTs within the optimal sparsity range while still capitalizing on the compression gains offered by random connectivity initialization. We explore the performance of SLTs within a frozen source network, i.e., a random network that (in addition to randomly pruned parameters) has randomly chosen parameters forced to be permanent part (i.e., never pruned) of the SLT—locked parameters (see Figure 2, right). Since the random locking pattern can be reconstructed from seed in the same way as the random pre-pruning pattern, our freezing method allows us to compress SLTs further. Additionally, locking allows us to extend the benefits of pre-pruning to the scenarios where the SLT sparsity should not be too high, e.g., as found in graph neural networks (Huang et al., 2022; Yan et al., 2023).
This section first describes the construction of frozen networks and the freezing pattern encoding for model compression. Then, we perform a preliminary experiment of the optimal settings for the proposed method in preparation for the evaluation experiments in Section 4.
3.1 Partial Freezing for Enhanced SLT Compression
Figure 3: Pre-pruning and locking set the bounds of the SLT sparsity that can be found. These optimal bounds are investigated in Section 3.2.
In our method, since frozen regions in the source network are completely random and thus can be reconstructed from seed, we need to memorize only the optimized supermask region. The total amount of frozen parameters is determined by a global freezing ratio $F_{r}$, which is the sum of the respective pre-pruning ratio $P_{r}$ and locking ratio $L_{r}$. Therefore, we optimize the $1\!-\!F_{r}$ of the parameters (i.e., non-frozen parameters) of the original dense network. As explained previously, it is not possible to search for an SLT with a lower sparsity (denser) than the frozen source network. Consequently, as shown in Figure 3, the pre-pruning ratio $P_{r}$ sets the lower bound of the sparsity of the SLTs that can be found. On the other hand, it is not possible to search for an SLT that prunes more parameters than those available, so the locking ratio $L_{r}$ sets the upper bound of the sparsity of the SLTs that can be found. (Specifically, the upper bound is given by $1\!-\!L_{r}$.) Thus, these ratios allow us to freely control the proportion of pre-pruning and locking ratios and memory size reduction ratio for the desired SLT sparsity.
Frozen network construction:
We assume that the $l$-th layer of network with depth $L$ has weights $\bm{w}^{(l)}\in\mathbb{R}^{n^{(l)}}$. The number of frozen parameters is determined by the layer-wise freezing ratios $p_{f}^{(1)},...,p_{f}^{(L)}$, where each $p_{f}^{(l)}$ is defined as the sum of the layer-wise pre-pruning ratio $p_{p}^{(l)}$ and the layer-wise locking ratio $p_{l}^{(l)}$, and $p_{f}^{(l)},p_{p}^{(l)},p_{l}^{(l)}\in[0,1]$.
To prune $p_{f}^{(l)}n^{(l)}$ parameters and lock $p_{l}^{(l)}n^{(l)}$ parameters, we generate two random masks, a pre-pruning mask ${\bm{m}^{(l)}_{p}\in\{0,1\}^{n^{(l)}}}$ and a locking mask ${\bm{m}^{(l)}_{l}\in\{0,1\}^{n^{(l)}}}$, so that they satisfy ${||1-\bm{m}^{(l)}_{p}||=p_{p}^{(l)}n^{(l)}}$, ${||\bm{m}^{(l)}_{l}||=p_{l}^{(l)}n^{(l)}}$, and ${(1-\bm{m}^{(l)}_{p})\cdot\bm{m}^{(l)}_{l}=0}$ (i.e., we require pre-pruning and locking to be implemented without overlap). The layer-wise weights frozen with these masks are calculated as ${\bm{w}^{(l)}_{f}:=(\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})+\bm{m}_{l}^{(l)})\odot\bm{w}^{(l)}}$. These masks are fixed during training, and we search for SLTs only in the parts where ${\bm{m}_{p}^{(l)}\odot(1-\bm{m}_{l}^{(l)})}$ is one.
Setting the layer-wise ratios:
Our method considers the following two existing strategies for determining the layer-wise pre-pruning ratio from the desired global pre-pruning ratio of the network:
  • Erdős-Rényi-Kernel (ERK): The pre-pruning ratio of layer $l$ is proportional to the scale ${(C_{in}^{(l)}+C_{out}^{(l)}+k_{h}^{(l)}+k_{w}^{(l)})/(C_{in}^{(l)}\cdot C_{out}^{(l)}\cdot k_{h}^{(l)}\cdot k_{w}^{(l)})}$, where $C_{in}^{(l)}$, $C_{out}^{(l)}$, $k_{h}^{(l)}$, and $k_{w}^{(l)}$ denote input channels, output channels, kernel height, and kernel width of the layer $l$, respectively (Evci et al., 2020).
  • Edge Per Layer (EPL): Each layer’s pre-pruning ratio is set so that they all have the same number of remaining weights (Price & Tanner, 2021; Gadhikar et al., 2023).
The same strategy is employed to determine the layer-wise pre-pruning and freezing ratios from their respective global ratios, and then the locking ratios are calculated as the difference between the corresponding freezing and pre-pruning ratios.
Freezing pattern encoding for model compression:
The freezing pattern can be encoded during inference as a ternary mask—a freezing mask—that indicates whether a parameter is pruned, locked, or part of the supermask. For example, encoding pre-pruning as $-1$, locking as $+1$, and supermask inclusion as 00, the layer-wise freezing mask can be encoded as $\bm{m}^{(l)}_{l}+(\bm{m}^{(l)}_{p}-1)\in\{-1,0,1\}^{n^{(l)}}$. Since this freezing mask is also random and fixed, it can be regenerated from its seed and ratios, similarly to the random weights. Furthermore, the supermask size is reduced by excluding from it the frozen parameters, so the SLTs found by this method can be compressed in inference to an even smaller memory size than those produced by the existing SLT literature (Hirose et al., 2022; Okoshi et al., 2022; López García-Arias et al., 2023), reducing costly off-chip memory access on specialized neural inference accelerators (Hirose et al., 2022; Chen et al., 2023), and thus offering an opportunity to perform even faster and more energy-efficient inference processing.
3.2 Optimal Pruning:Locking Proportion for Freezing
Here, we perform a preliminary investigation of the optimal pruning:locking proportion for each given desired SLT sparsity by varying the proportion with a fixed freezing ratio of the network. Figure 4 explores different configurations of a $80\%$ freezing ratio—the situation where the supermask memory size is $20\%$ of the SLT in the dense source network—on a Conv6 network and compares the performance of the found SLTs on CIFAR-1010{10}10. As expected, the best-performing SLT of each prune:lock configuration is found at the center of the non-frozen region, where the number of candidate subnetworks is maximized. In other words, when the non-frozen region accounts for $S\in[0,1]$ of the entire source network, the optimal SLT sparsity is $k=P_{r}+S/2$, where $P_{r}=\sum_{l}p^{(l)}_{p}n^{(l)}/\sum_{k}n^{(k)}$ is the global pre-pruning ratio of the network. Conversely, for a given freezing ratio $F_{r}$ of the network and a desired SLT sparsity $k$, the optimal position of the frozen region is set by the pre-pruning ratio $P_{r}=k-(1-F_{r})/2$ and the corresponding locking ratio $L_{r}=F_{r}-P_{r}$. In the cases where this would result in $P_{r}\!<\!0$ or $L_{r}\!<\!0$, we choose a best-effort approach that keeps the desired freezing ratio and sets the bounds to $P_{r}\!=\!0$ or $L_{r}\!=\!0$, respectively.
Additionally, Figure 4 compares the two strategies for setting ratios considered in Section 3.1, showing that EPL outperforms ERK in all cases. Consequently, hereafter, the proposed method sets the global ratios in order to position the frozen region center as close as possible to the desired SLT sparsity, and then sets the layer-wise ratios using EPL.